OpenStudy (anonymous):

Simplify the expression –3(x + 3)2 – 3 + 3x. What is the simplified expression in standard form? –3x2 – 18x – 27 –3 – 15x – 30 –3 + 3x + 6 –3 + 3x – 30

OpenStudy (alexandervonhumboldt2):

here

OpenStudy (alexandervonhumboldt2):

-3x^2-18x-27-3+3x=-3x^2-15x-30

OpenStudy (anonymous):

So A ?

OpenStudy (alexandervonhumboldt2):

hmmno there is no answer @AriPotta

OpenStudy (alexandervonhumboldt2):

maybe i did a mistake

OpenStudy (alexandervonhumboldt2):

@dtuniyants

OpenStudy (anonymous):

just a sec

OpenStudy (alexandervonhumboldt2):

–3(x + 3)2 – 3 + 3x=-3* (x^2+6x+9)-3+3x=-3x^2-18x-27-3+3x=-3x^2-15x-30

OpenStudy (alexandervonhumboldt2):

nothing is right

OpenStudy (anonymous):

-3(x+3)^2-3+3x=-3(x^2+6x+9)-3+3x=-3x^2-18x-27-3+3x=-3x^2-15x-30

OpenStudy (alexandervonhumboldt2):

@satellite73

OpenStudy (anonymous):

I think its B. It is most similar to the real answer. They might have mistaken and forgot to write x^2 with -3.